OpenStudy (anonymous):

Convert the following complex number into its polar representation: -4i

OpenStudy (zzr0ck3r):

what will the radius be?

OpenStudy (zzr0ck3r):

think about the point (0,4) what angle goes to this point?

OpenStudy (anonymous):

not sure :/

OpenStudy (jdoe0001):

the complex standard form is "a+bi" you have pretty much as zzr0ck3r already suggested 0-4i so 0 -4i a b (a, b) elements

OpenStudy (anonymous):

a

OpenStudy (anonymous):

b

OpenStudy (anonymous):

c

OpenStudy (zzr0ck3r):

um those are not polar....

OpenStudy (anonymous):

d

OpenStudy (jdoe0001):

so you'd want to convert it to the polar form of \(\large r = (cos(\theta) + isin(\theta)\)

OpenStudy (anonymous):

these are my answer choices :/

OpenStudy (zzr0ck3r):

polar = (radius,angle)

OpenStudy (jdoe0001):

so, bear in mind that \(r = a^2 + b^2\)

OpenStudy (jdoe0001):

woops darn

OpenStudy (jdoe0001):

\(r^2 = a^2 + b^2\)

OpenStudy (anonymous):

its C?

OpenStudy (jdoe0001):

dunno

OpenStudy (anonymous):

so you dont know ?

OpenStudy (jdoe0001):

well, we haven't gotten the value for "r"

OpenStudy (anonymous):

how do you do that